OpenStudy (anonymous):

x+2y+z=10 2x-y+3z=-5 2x-3y-5z=27 Please solve step by step please, Thanks!

OpenStudy (mathmate):

What methods have you been working on, elimination, Cramers, row reduction or inverse?

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

sorry was away

OpenStudy (mathmate):

x+2y+z=10 ...(1) 2x-y+3z=-5 ...(2) 2x-3y-5z=27...(3) 2(1)-2: 5y-z = 25 ...(4) (2)-(3): 2y+8z=-32 ...(5) 2(4)-5(5): -42z=210 => z=-5 ...(6) substitute (6) into (4) 5y-(-5)=25 y=(25-5)/5=4 ....(7) Substitute (6) and (7) into (1) x+2(4)+(-5)=10 => x=7 So (x,y,z)=(7,4,-5)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Omg i can't thank you enough

OpenStudy (mathmate):

You're welcome!